# Measuring Social Norms of Large Language Models

### v2

@@ Appendix D Summary of Skills
Subject: Language Arts Skill: to-be-use-the-correct-form Description: Complete the sentence with the correct **present tense** form of **to be**. They ___ upstairs in Sofia’s bedroom. Choices: [(a) are, (b) am] Answer index: 0
Subject: Language Arts Skill: read-about-famous-places Description: Review the third paragraph. The Golden Gate Bridge is one of the most famous bridges in the world. Its tall frame, two large towers, and orange color make the bridge easy to recognize. Yet as the Golden Gate Bridge was being built, people disagreed about what the bridge should look like. People were worried that San Francisco’s thick fog would make the bridge hard to see. The U.S. Navy wanted the bridge to stand out so ships and airplanes wouldn’t collide with it. They wanted the bridge to have wide black and yellow stripes, like a bumblebee! Irving Morrow was one of the bridge designers. He did not like the idea of a striped bridge. Morrow did not want a plain black or gray bridge either, though. He thought a dark bridge would make the Golden Gate seem smaller than it was. When the bridge’s parts were delivered to San Francisco, they had an orange color. The bridge workers were supposed to paint a new color over the orange. But Morrow was still unsure of what color that would be. He looked at the orange bridge parts and decided that the color would be perfect! Morrow thought the orange would stand out well against the ocean and the sky. Finally, the difficult decision had been made. A red-orange Golden Gate Bridge was bright enough for ships and planes to see. The color made the bridge appear grand and tall. Over the years, fresh paint has been needed. But the color remains the same orange that was supposed to be painted over many years ago. Based on the text, why did Irving Morrow think orange would be a good color for the Golden Gate Bridge? Choices: [(a) The parts were already orange, so no extra work would be needed., (b) The color would make the bridge stand out from the sky and ocean., (c) Using orange paint on orange parts would be cheaper.] Answer index: 1
Subject: Language Arts Skill: does-the-adverb-tell-you-how-when-or-where Description: Look at the adverb in bold. Does it tell you "how" or "where"? Mr. Olson lives **upstairs**. Choices: [(a) how, (b) where] Answer index: 1
Subject: Language Arts Skill: which-word-is-a-noun Description: Is the word in bold a noun? The captain turned the ship around and **sailed** north. Choices: [(a) yes, (b) no] Answer index: 1
@@
Subject: Language Arts Skill: identify-the-irregular-past-tense-ii Description: Which sentence is in the past tense? Choices: [(a) Susan **feeds** her bird., (b) Susan **fed** her bird.] Answer index: 1
Subject: Language Arts Skill: which-sentence-is-in-the-regular-past-tense Description: Which sentence is in the past tense? Choices: [(a) Cassie **acted** happy., (b) We **discover** hidden treasures.] Answer index: 0
@@
Subject: Language Arts Skill: which-sentence-uses-an-antonym Description: Read the sentence. Select the sentence that uses an antonym in place of the word in bold. Bella’s dad has always been **short**. Choices: [(a) Bella’s dad has always been **tall**., (b) Bella’s dad has always been **shy**.] Answer index: 0
Subject: Language Arts Skill: use-action-verbs Description: Complete the sentence with the action verb. Mrs. Carlson ___ nuts into the cookie batter. Choices: [(a) mixes, (b) butter] Answer index: 0
Subject: Language Arts Skill: pronoun-verb-agreement Description: Complete the sentence with the best **verb**. He ___ on television. Choices: [(a) appears, (b) appear] Answer index: 0
Subject: Language Arts Skill: make-predictions-about-a-story Description: Read this passage from a story. Anna and Andrew had just moved in to their new house. "Let’s explore," Anna said. They opened all the closets in the bedrooms. Then they found the door to the attic. Their dad noticed them opening the door. "Go ahead," he said. "It’s full of old junk." In the attic, Anna and Andrew saw an old dresser and some boards. Then they noticed a very large metal box. They walked over to look at it more closely. Based on the passage, which is more likely to happen next? Choices: [(a) Anna and Andrew open the metal box., (b) Anna and Andrew open the dresser drawers.] Answer index: 0
Subject: Language Arts Skill: greetings-and-closings-of-letters Description: Which closing is correct for a letter? Choices: [(a) Thanks, Jim, (b) thanks, Jim] Answer index: 0
Subject: Language Arts Skill: identify-base-words-prefixes-and-suffixes Description: What is the base word in **dampness**? Choices: [(a) damp, (b) ness] Answer index: 0
Subject: Language Arts Skill: which-sentence-has-the-same-meaning Description: Which sentence has the same meaning? Mrs. Duncan **closed** her bedroom door and went to bed. Choices: [(a) Mrs. Duncan **shut** her bedroom door and went to bed., (b) Mrs. Duncan **painted** her bedroom door and went to bed.] Answer index: 0
Subject: Language Arts Skill: identify-the-authors-purpose-passages Description: Read the text below. Vote for Bob Mueller for city council. He’ll stand up for the needs of our town’s hardworking citizens. What is the text’s most likely purpose? Choices: [(a) to persuade, (b) to inform, (c) to entertain] Answer index: 0
@@
Subject: Language Arts Skill: choose-topic-sentences-for-narrative-paragraphs Description: Read the paragraph. Then, add the better topic sentence. ___ First, my parents took me to the store to buy some seeds. Then, we took the seeds home, and I planted them in my backyard. I watered them every day, and soon the plants started to grow. After several weeks, I started to see little vegetables pop up. By the end of the summer, I had a garden full of green peppers, beans, corn, and carrots. Choices: [(a) Last spring, I planted an amazing vegetable garden., (b) Fresh vegetables are my favorite topping to put on pizza.] Answer index: 0
@@
Subject: Language Arts Skill: choose-the-text-that-matches-the-writers-purpose Description: You want to **inform** someone **about blueberries**. "Inform" means "teach someone about something". What should you write? Choices: [(a) Sarah couldn’t believe her eyes. On the table were plates full of blueberry pancakes, blueberry muffins, and even blueberry sauces., (b) Order our blueberry pancakes today! They’re always hot and delicious., (c) As the blueberry fruit ripens, it turns from green to purple to dark blue.] Answer index: 2
Subject: Language Arts Skill: analyze-passages-from-narrative-of-the-life-of-frederick-douglass-part-2 Description: Review the passage. I thought the matter over during the next day, Sunday, and finally resolved upon the third day of September, as the day upon which I would make a second attempt to secure my freedom. I now had three weeks during which to prepare for my journey. From Frederick Douglass, Narrative of the Life of Frederick Douglass Look at the text in **bold**. What hidden reality is shown through this text? Choices: [(a) It shows that although Douglass’s friends behaved kindly, they secretly envied him., (b) It shows that although the shipyard was calm, Douglass experienced tension at home., (c) It shows that although the city felt peaceful, enslaved people were preparing to rebel., (d) It shows that although Douglass’s plan was working, he still had worries and doubts.] Answer index: 3
@@
Subject: Language Arts Skill: are-the-modifiers-used-correctly Description: Which sentence is correct? Choices: [(a) **Stuck in traffic**, Hakim grew drowsy and time slowed to a standstill; eventually, he decided to pull over and take a nap., (b) **Stuck in traffic**, time slowed to a standstill and Hakim grew drowsy; eventually, he decided to pull over and take a nap.] Answer index: 0
Subject: Language Arts Skill: recall-the-source-of-an-allusion Description: What is the source of the allusion in the sentence below? The headline read: "Smart Phones Becoming **Big Brother**." Choices: [(a) literature, (b) a song] Answer index: 0
@@
Subject: Language Arts Skill: transitions-with-conjunctive-adverbs Description: Complete the text with the better conjunctive adverb. In China, mountainous areas are being flattened to make more room for new construction; ___, in the United States, mountaintops are being removed so that miners can access coal reserves. Choices: [(a) therefore, (b) likewise] Answer index: 1
@@
Subject: Language Arts Skill: determine-the-authors-point-of-view Description: Read the text. When I decided it was time for a new hobby, I signed up for a "Clowning for Fun and Profit" class at my local arts center. I cheerily skipped through the door on the first day of class, but my heart sank as soon as I spotted the instructor, Bobo. Bobo turned out to be a sad-faced clown who refused to speak. It was clear he planned to teach entirely in pantomime. As he silently motioned for me to find my seat, I resigned myself to a bad experience. So imagine my surprise when this crazy class turned out to be invaluable to me! Although I didn’t hear a single spoken word for five hours, Bobo somehow managed to express his ideas clearly and in an engaging way. By the end of the first session, he had expertly taught me how to make balloon animals, apply makeup, juggle beanbags, and develop my own personal clown character. As I eagerly await the second session, "All About Unicycles," I will begin clowning around as a volunteer at a local children’s hospital. Which sentence best summarizes the author’s point of view about the clown class? Choices: [(a) Bobo uses eccentric teaching methods to efficiently teach this course, making him an excellent clowning instructor., (b) Bobo’s sad demeanor and poor pantomime skills make it difficult for his students to learn effective clowning techniques.] Answer index: 0
@@
Subject: Language Arts Skill: compare-two-texts-with-different-genres Description: Read the following text from a diary. This excerpt is about life for Jews in Europe at the beginning of World War II. After May 1940 good times rapidly fled: first the war, then the capitulation, followed by the arrival of the Germans, which is when the sufferings of us Jews really began. Anti-Jewish decrees followed each other in quick succession. Jews must wear a yellow star, Jews must hand in their bicycles, Jews are banned from trams and are forbidden to drive. Jews are only allowed to do their shopping between three and five o’clock and then only in shops which bear the placard "Jewish shop." Jews must be indoors by eight o’clock and cannot even sit in their own gardens after that hour. Jews are forbidden to visit theaters, cinemas, and other places of entertainment. From Anne Frank, "Anne Frank: The Diary of a Young Girl". Copyright 1952 by Otto H. Frank Based on this excerpt, what can you conclude about life for Jews in Europe at the beginning of World War II? Choices: [(a) Jewish people could no longer afford to buy theater tickets., (b) Jewish people could not associate with one another., (c) Jewish people had many of their rights taken away.] Answer index: 2
@@
Subject: Language Arts Skill: read-and-understand-informational-passages Description: Review the **seventh paragraph**. title: Meditating in School This is where the need for good, sound research comes in. Past studies showing that mindfulness helps medical patients are not enough. To really determine whether mindfulness helps improve grades, research needs to focus specifically on mindfulness in schools—and what impact, if any, it has. That way, educators can make informed decisions about where to put their resources. Until then, whether mindfulness is included in classroom programming will remain the choice of individual teachers. Based on the text, what still needs to be determined about mindfulness in schools? Choices: [(a) whether meditation affects different students in different ways, (b) how to raise the funds needed to teach meditation in schools, (c) whether mindfulness meditation has a positive impact in a school setting] Answer index: 2
@@
Subject: Language Arts Skill: choose-evidence-to-support-a-claim Description: Read the claim below. Gardeners should never use pesticides on their plants. Select the piece of evidence that best supports this claim. Choices: [(a) Organic pesticides tend to cost more than conventional ones, and they may be less effective., (b) Pets and children often play in and around gardens, where they can be exposed to pesticides.] Answer index: 1
@@
Subject: Language Arts Skill: analyze-passages-from-the-giver-part-1 Description: Reread the following passage from The Giver. In this excerpt, Jonas, the main character, recalls a time when he saw an unidentified aircraft. Frightened was the way [Jonas] had felt a year ago when an unidentified aircraft had overflown the community twice. He had seen it both times. Squinting toward the sky, he had seen the sleek jet, almost a blur at its high speed, go past, and a second later heard the blast of sound that followed. Then one more time, a moment later, from the opposite direction, the same plane. From Lois Lowry, The Giver. Copyright 1993 by Lois Lowry Look at the text in **bold**. How does this paragraph contribute to the meaning or tone of the passage? Choices: [(a) It creates an unsettling tone., (b) It creates a ridiculous tone., (c) It creates a playful tone., (d) It creates a serene tone.] Answer index: 0
Subject: Language Arts Skill: use-personification Description: Complete the sentence so that it uses personification. The sorceress peered into the mirror, noticing how the amulet ___ her neck. Choices: [(a) barely clasped around, (b) gently embraced] Answer index: 1
Subject: Language Arts Skill: identify-plagiarism Description: Compare the student text with the source. Source: Carolyn Cinami DeCristofano, "A Black Hole Is Not a Hole", page 5. Published by Charlesbridge, 2012: Nothing can out-tug a black hole. No army of tow trucks, no convoy of supersized earth haulers, no fleet of giant rocket engines. Student text: The pull of a black hole is so strong that nothing can pull stronger, not even a "fleet of giant rocket engines" (DeCristofano 5). Is the student text plagiarized? Choices: [(a) No, it is not plagiarized., (b) Yes, because it fails to use quotation marks.] Answer index: 0
Subject: Language Arts Skill: identify-counterclaims Description: Consider this claim: Poems that rhyme are better than poems that do not rhyme. Now consider how someone might argue against this. Which sentence presents a counterclaim to the above claim? Choices: [(a) "Beowulf" is one of the oldest and longest poems in existence., (b) Poetry that doesn’t rhyme allows for more freedom of expression.] Answer index: 1
@@
Subject: Language Arts Skill: trace-an-argument Description: Read the text. title: Early to Bed and Early to Rise Whether you’re naturally a night owl or an early bird is mostly based on your body’s chronotype, which Merriam-Webster’s online dictionary defines as "the internal circadian rhythm or body clock of an individual that influences the cycle of sleep and activity in a 24-hour period." In short, we are born wired one way or the other. However, it is possible to change your habits. And given the benefits of being an early riser, it may be worth the effort. What is the author’s main claim or argument? Choices: [(a) There are just as many benefits to being a night owl as an early riser., (b) It’s important to change how society perceives night owls., (c) It is better to be an early riser than a night owl.] Answer index: 2
Subject: Language Arts Skill: identify-active-and-passive-voice Description: Which sentence is in passive voice? Choices: [(a) On April 18, 1906, an earthquake and its fiery aftermath devastated the city of San Francisco., (b) On April 18, 1906, San Francisco was devastated by an earthquake and its fiery aftermath.] Answer index: 1
@@
Subject: Language Arts Skill: is-the-pronoun-reflexive-or-intensive Description: Is the pronoun in bold reflexive or intensive? Bryce uses an elaborate system of notes to remind **himself** of what he has left to do. Choices: [(a) intensive, (b) reflexive] Answer index: 1
Subject: Language Arts Skill: interpret-figures-of-speech Description: What does the **metaphor** in this text suggest? When Bobby lost his job, he was devastated. The only light in the sea of darkness was the prospect of pursuing a new career. Choices: [(a) There was a benefit to Bobby’s job loss., (b) Having to pursue a new career was the worst part of Bobby’s job loss.] Answer index: 0
Subject: Language Arts Skill: identify-the-verb-mood Description: Which verb mood is used in the **bold** part of the sentence? To help maintain your energy and focus throughout the day, **eat a healthy breakfast**. Choices: [(a) imperative, (b) subjunctive] Answer index: 0
Subject: Language Arts Skill: classify-figures-of-speech Description: Which figure of speech is used in this text? The water made a sound **like kittens lapping**. —Marjorie Kinnan Rawlings, "The Yearling" Choices: [(a) metaphor, (b) simile] Answer index: 1
@@
Subject: Language Arts Skill: vocabulary-review-trace-an-argument Description: Look at the word **entrenched** as used in the passage "The Most Important Meal of the Day?" An Australian study published in 2019 collected health data for over twenty-eight years from multiple countries. The results showed few differences between breakfast skippers and breakfast eaters. The study concluded that "breakfast is not the most important time of the day to eat, even though that belief is really **entrenched** in our society and around the world." What does **entrenched** mean? Choices: [(a) highly debatable, (b) quite rare, (c) firmly established] Answer index: 2
@@
Subject: Language Arts Skill: analyze-passages-from-the-outsiders-part-2 Description: Review the passage. Tim Shepard and company were already waiting when we arrived at the vacant lot, along with a gang from Brumly, one of the suburbs. Tim was a lean, catlike eighteen-year-old who looked like the model JD you see in movies and magazines. He had the right curly black hair, smoldering dark eyes, and a long scar from temple to chin where a tramp had belted him with a broken pop bottle. He had a tough, hard look to him, and his nose had been broken twice. Like Dally’s, his smile was grim and bitter. He was one of those who enjoy being a hood. The rest of his bunch were the same way. The boys from Brumly, too. Young hoods—who would grow up to be old hoods. I’d never thought about it before, but they’d just get worse as they got older, not better. I looked at Darry. He wasn’t going to be any hood when he got old. He was going to get somewhere. Living the way we do would only make him more determined to get somewhere. That’s why he’s better than the rest of us, I thought. He’s going somewhere. And I was going to be like him. I wasn’t going to live in a lousy neighborhood all my life. From S. E. Hinton, The Outsiders. Copyright 1967 by S. E. Hinton Look at the text in **bold**. What does this figurative language suggest? Choices: [(a) Tim fights for survival every day., (b) Tim has always been cruel to Ponyboy., (c) Tim often does not have enough to eat., (d) Tim attempts to stay hidden at all costs.] Answer index: 0

### v3

@@ Appendix D Summary of Skills
Description: Complete the sentence with the correct **present tense** form of **to be**.
They ___ upstairs in Sofia’s bedroom.
Choices: [(a) are, (b) am]
Answer index: 0
Subject: Language Arts
Skill: read-about-famous-places
Description: Review the third paragraph.
The Golden Gate Bridge is one of the most famous bridges in the world. Its tall frame, two large towers, and orange color make the bridge easy to recognize. Yet as the Golden Gate Bridge was being built, people disagreed about what the bridge should look like.
People were worried that San Francisco’s thick fog would make the bridge hard to see. The U.S. Navy wanted the bridge to stand out so ships and airplanes wouldn’t collide with it. They wanted the bridge to have wide black and yellow stripes, like a bumblebee! Irving Morrow was one of the bridge designers. He did not like the idea of a striped bridge. Morrow did not want a plain black or gray bridge either, though. He thought a dark bridge would make the Golden Gate seem smaller than it was.
When the bridge’s parts were delivered to San Francisco, they had an orange color. The bridge workers were supposed to paint a new color over the orange. But Morrow was still unsure of what color that would be. He looked at the orange bridge parts and decided that the color would be perfect! Morrow thought the orange would stand out well against the ocean and the sky.
Finally, the difficult decision had been made. A red-orange Golden Gate Bridge was bright enough for ships and planes to see. The color made the bridge appear grand and tall. Over the years, fresh paint has been needed. But the color remains the same orange that was supposed to be painted over many years ago.
Based on the text, why did Irving Morrow think orange would be a good color for the Golden Gate Bridge?
Choices: [(a) The parts were already orange, so no extra work would be needed., (b) The color would make the bridge stand out from the sky and ocean., (c) Using orange paint on orange parts would be cheaper.]
Answer index: 1
Subject: Language Arts
Skill: does-the-adverb-tell-you-how-when-or-where
Description: Look at the adverb in bold. Does it tell you "how" or "where"?
Mr. Olson lives **upstairs**.
Choices: [(a) how, (b) where]
Answer index: 1
Subject: Language Arts
Skill: which-word-is-a-noun
Description: Is the word in bold a noun?
The captain turned the ship around and **sailed** north.
Choices: [(a) yes, (b) no]
Answer index: 1
@@
Subject: Language Arts
Skill: identify-the-irregular-past-tense-ii
Description: Which sentence is in the past tense?
Choices: [(a) Susan **feeds** her bird., (b) Susan **fed** her bird.]
Answer index: 1
Subject: Language Arts
Skill: which-sentence-is-in-the-regular-past-tense
Description: Which sentence is in the past tense?
Choices: [(a) Cassie **acted** happy., (b) We **discover** hidden treasures.]
Answer index: 0
Subject: Language Arts
Skill: which-sentence-uses-an-antonym
@@
Description: Read the sentence. Select the sentence that uses an antonym in place of the word in bold.
Bella’s dad has always been **short**.
Choices: [(a) Bella’s dad has always been **tall**., (b) Bella’s dad has always been **shy**.]
Answer index: 0
Subject: Language Arts
Skill: use-action-verbs
Description: Complete the sentence with the action verb.
Mrs. Carlson ___ nuts into the cookie batter.
Choices: [(a) mixes, (b) butter]
Answer index: 0
Subject: Language Arts
Skill: pronoun-verb-agreement
Description: Complete the sentence with the best **verb**.
He ___ on television.
Choices: [(a) appears, (b) appear]
Answer index: 0
Subject: Language Arts
Skill: make-predictions-about-a-story
Description: Read this passage from a story.
Anna and Andrew had just moved in to their new house. "Let’s explore," Anna said. They opened all the closets in the bedrooms. Then they found the door to the attic. Their dad noticed them opening the door. "Go ahead," he said. "It’s full of old junk."
In the attic, Anna and Andrew saw an old dresser and some boards. Then they noticed a very large metal box. They walked over to look at it more closely.
Based on the passage, which is more likely to happen next?
Choices: [(a) Anna and Andrew open the metal box., (b) Anna and Andrew open the dresser drawers.]
Answer index: 0
Subject: Language Arts
Skill: greetings-and-closings-of-letters
Description: Which closing is correct for a letter?
Choices: [(a) Thanks, Jim, (b) thanks, Jim]
Answer index: 0
Subject: Language Arts
Skill: identify-base-words-prefixes-and-suffixes
Description: What is the base word in **dampness**?
Choices: [(a) damp, (b) ness]
Answer index: 0
Subject: Language Arts
Skill: which-sentence-has-the-same-meaning
Description: Which sentence has the same meaning?
Mrs. Duncan **closed** her bedroom door and went to bed.
Choices: [(a) Mrs. Duncan **shut** her bedroom door and went to bed., (b) Mrs. Duncan **painted** her bedroom door and went to bed.]
Answer index: 0
Subject: Language Arts
Skill: identify-the-authors-purpose-passages
Description: Read the text below.
Vote for Bob Mueller for city council. He’ll stand up for the needs of our town’s hardworking citizens.
What is the text’s most likely purpose?
Choices: [(a) to persuade, (b) to inform, (c) to entertain]
Answer index: 0
Subject: Language Arts
Skill: choose-topic-sentences-for-narrative-paragraphs
Description: Read the paragraph. Then, add the better topic sentence.
@@
___ First, my parents took me to the store to buy some seeds. Then, we took the seeds home, and I planted them in my backyard. I watered them every day, and soon the plants started to grow. After several weeks, I started to see little vegetables pop up. By the end of the summer, I had a garden full of green peppers, beans, corn, and carrots.
Choices: [(a) Last spring, I planted an amazing vegetable garden., (b) Fresh vegetables are my favorite topping to put on pizza.]
Answer index: 0
Subject: Language Arts
Skill: choose-the-text-that-matches-the-writers-purpose
Description: You want to **inform** someone **about blueberries**. "Inform" means "teach someone about something". What should you write?
@@
Choices: [(a) Sarah couldn’t believe her eyes. On the table were plates full of blueberry pancakes, blueberry muffins, and even blueberry sauces., (b) Order our blueberry pancakes today! They’re always hot and delicious., (c) As the blueberry fruit ripens, it turns from green to purple to dark blue.]
Answer index: 2
Subject: Language Arts
Skill: analyze-passages-from-narrative-of-the-life-of-frederick-douglass-part-2
Description: Review the passage.
I thought the matter over during the next day, Sunday, and finally resolved upon the third day of September, as the day upon which I would make a second attempt to secure my freedom. I now had three weeks during which to prepare for my journey.
From Frederick Douglass, Narrative of the Life of Frederick Douglass
Look at the text in **bold**. What hidden reality is shown through this text?
Choices: [(a) It shows that although Douglass’s friends behaved kindly, they secretly envied him., (b) It shows that although the shipyard was calm, Douglass experienced tension at home., (c) It shows that although the city felt peaceful, enslaved people were preparing to rebel., (d) It shows that although Douglass’s plan was working, he still had worries and doubts.]
Answer index: 3
Subject: Language Arts
Skill: are-the-modifiers-used-correctly
Description: Which sentence is correct?
@@
Choices: [(a) **Stuck in traffic**, Hakim grew drowsy and time slowed to a standstill; eventually, he decided to pull over and take a nap., (b) **Stuck in traffic**, time slowed to a standstill and Hakim grew drowsy; eventually, he decided to pull over and take a nap.]
Answer index: 0
Subject: Language Arts
Skill: recall-the-source-of-an-allusion
Description: What is the source of the allusion in the sentence below?
The headline read: "Smart Phones Becoming **Big Brother**."
Choices: [(a) literature, (b) a song]
Answer index: 0
Subject: Language Arts
Skill: transitions-with-conjunctive-adverbs
Description: Complete the text with the better conjunctive adverb.
@@
In China, mountainous areas are being flattened to make more room for new construction; ___, in the United States, mountaintops are being removed so that miners can access coal reserves.
Choices: [(a) therefore, (b) likewise]
Answer index: 1
Subject: Language Arts
Skill: determine-the-authors-point-of-view
Description: Read the text.
When I decided it was time for a new hobby, I signed up for a "Clowning for Fun and Profit" class at my local arts center. I cheerily skipped through the door on the first day of class, but my heart sank as soon as I spotted the instructor, Bobo. Bobo turned out to be a sad-faced clown who refused to speak. It was clear he planned to teach entirely in pantomime. As he silently motioned for me to find my seat, I resigned myself to a bad experience.
@@
So imagine my surprise when this crazy class turned out to be invaluable to me! Although I didn’t hear a single spoken word for five hours, Bobo somehow managed to express his ideas clearly and in an engaging way. By the end of the first session, he had expertly taught me how to make balloon animals, apply makeup, juggle beanbags, and develop my own personal clown character. As I eagerly await the second session, "All About Unicycles," I will begin clowning around as a volunteer at a local children’s hospital.
Which sentence best summarizes the author’s point of view about the clown class?
Choices: [(a) Bobo uses eccentric teaching methods to efficiently teach this course, making him an excellent clowning instructor., (b) Bobo’s sad demeanor and poor pantomime skills make it difficult for his students to learn effective clowning techniques.]
Answer index: 0
Subject: Language Arts
Skill: compare-two-texts-with-different-genres
Description: Read the following text from a diary. This excerpt is about life for Jews in Europe at the beginning of World War II.
@@
After May 1940 good times rapidly fled: first the war, then the capitulation, followed by the arrival of the Germans, which is when the sufferings of us Jews really began. Anti-Jewish decrees followed each other in quick succession. Jews must wear a yellow star, Jews must hand in their bicycles, Jews are banned from trams and are forbidden to drive. Jews are only allowed to do their shopping between three and five o’clock and then only in shops which bear the placard "Jewish shop." Jews must be indoors by eight o’clock and cannot even sit in their own gardens after that hour. Jews are forbidden to visit theaters, cinemas, and other places of entertainment.
From Anne Frank, "Anne Frank: The Diary of a Young Girl". Copyright 1952 by Otto H. Frank
Based on this excerpt, what can you conclude about life for Jews in Europe at the beginning of World War II?
Choices: [(a) Jewish people could no longer afford to buy theater tickets., (b) Jewish people could not associate with one another., (c) Jewish people had many of their rights taken away.]
Answer index: 2
@@
Subject: Language Arts
Skill: read-and-understand-informational-passages
Description: Review the **seventh paragraph**.
title: Meditating in School
@@
This is where the need for good, sound research comes in. Past studies showing that mindfulness helps medical patients are not enough. To really determine whether mindfulness helps improve grades, research needs to focus specifically on mindfulness in schools—and what impact, if any, it has. That way, educators can make informed decisions about where to put their resources. Until then, whether mindfulness is included in classroom programming will remain the choice of individual teachers.
Based on the text, what still needs to be determined about mindfulness in schools?
Choices: [(a) whether meditation affects different students in different ways, (b) how to raise the funds needed to teach meditation in schools, (c) whether mindfulness meditation has a positive impact in a school setting]
Answer index: 2
Subject: Language Arts
Skill: choose-evidence-to-support-a-claim
Description: Read the claim below.
Gardeners should never use pesticides on their plants.
Select the piece of evidence that best supports this claim.
@@
Choices: [(a) Organic pesticides tend to cost more than conventional ones, and they may be less effective., (b) Pets and children often play in and around gardens, where they can be exposed to pesticides.]
Answer index: 1
Subject: Language Arts
Skill: analyze-passages-from-the-giver-part-1
Description: Reread the following passage from The Giver. In this excerpt, Jonas, the main character, recalls a time when he saw an unidentified aircraft.
@@
Frightened was the way [Jonas] had felt a year ago when an unidentified aircraft had overflown the community twice. He had seen it both times. Squinting toward the sky, he had seen the sleek jet, almost a blur at its high speed, go past, and a second later heard the blast of sound that followed. Then one more time, a moment later, from the opposite direction, the same plane.
From Lois Lowry, The Giver. Copyright 1993 by Lois Lowry
Look at the text in **bold**. How does this paragraph contribute to the meaning or tone of the passage?
Choices: [(a) It creates an unsettling tone., (b) It creates a ridiculous tone., (c) It creates a playful tone., (d) It creates a serene tone.]
Answer index: 0
Subject: Language Arts
Skill: use-personification
Description: Complete the sentence so that it uses personification.
The sorceress peered into the mirror, noticing how the amulet ___ her neck.
Choices: [(a) barely clasped around, (b) gently embraced]
Answer index: 1
Subject: Language Arts
Skill: identify-plagiarism
Description: Compare the student text with the source.
Source: Carolyn Cinami DeCristofano, "A Black Hole Is Not a Hole", page 5. Published by Charlesbridge, 2012:
Nothing can out-tug a black hole. No army of tow trucks, no convoy of supersized earth haulers, no fleet of giant rocket engines.
Student text:
The pull of a black hole is so strong that nothing can pull stronger, not even a "fleet of giant rocket engines" (DeCristofano 5).
Is the student text plagiarized?
Choices: [(a) No, it is not plagiarized., (b) Yes, because it fails to use quotation marks.]
Answer index: 0
Subject: Language Arts
Skill: identify-counterclaims
Description: Consider this claim:
Poems that rhyme are better than poems that do not rhyme.
Now consider how someone might argue against this. Which sentence presents a counterclaim to the above claim?
Choices: [(a) "Beowulf" is one of the oldest and longest poems in existence., (b) Poetry that doesn’t rhyme allows for more freedom of expression.]
Answer index: 1
Subject: Language Arts
Skill: trace-an-argument
Description: Read the text.
title: Early to Bed and Early to Rise
@@
Whether you’re naturally a night owl or an early bird is mostly based on your body’s chronotype, which Merriam-Webster’s online dictionary defines as "the internal circadian rhythm or body clock of an individual that influences the cycle of sleep and activity in a 24-hour period." In short, we are born wired one way or the other. However, it is possible to change your habits. And given the benefits of being an early riser, it may be worth the effort.
What is the author’s main claim or argument?
Choices: [(a) There are just as many benefits to being a night owl as an early riser., (b) It’s important to change how society perceives night owls., (c) It is better to be an early riser than a night owl.]
Answer index: 2
Subject: Language Arts
Skill: identify-active-and-passive-voice
Description: Which sentence is in passive voice?
Choices: [(a) On April 18, 1906, an earthquake and its fiery aftermath devastated the city of San Francisco., (b) On April 18, 1906, San Francisco was devastated by an earthquake and its fiery aftermath.]
Answer index: 1
Subject: Language Arts
Skill: is-the-pronoun-reflexive-or-intensive
Description: Is the pronoun in bold reflexive or intensive?
@@
Bryce uses an elaborate system of notes to remind **himself** of what he has left to do.
Choices: [(a) intensive, (b) reflexive]
Answer index: 1
Subject: Language Arts
Skill: interpret-figures-of-speech
Description: What does the **metaphor** in this text suggest?
When Bobby lost his job, he was devastated. The only light in the sea of darkness was the prospect of pursuing a new career.
Choices: [(a) There was a benefit to Bobby’s job loss., (b) Having to pursue a new career was the worst part of Bobby’s job loss.]
Answer index: 0
Subject: Language Arts
Skill: identify-the-verb-mood
Description: Which verb mood is used in the **bold** part of the sentence?
To help maintain your energy and focus throughout the day, **eat a healthy breakfast**.
Choices: [(a) imperative, (b) subjunctive]
Answer index: 0
Subject: Language Arts
Skill: classify-figures-of-speech
Description: Which figure of speech is used in this text?
The water made a sound **like kittens lapping**.
—Marjorie Kinnan Rawlings, "The Yearling"
Choices: [(a) metaphor, (b) simile]
Answer index: 1
Subject: Language Arts
Skill: vocabulary-review-trace-an-argument
Description: Look at the word **entrenched** as used in the passage "The Most Important Meal of the Day?"
@@
An Australian study published in 2019 collected health data for over twenty-eight years from multiple countries. The results showed few differences between breakfast skippers and breakfast eaters. The study concluded that "breakfast is not the most important time of the day to eat, even though that belief is really **entrenched** in our society and around the world."
What does **entrenched** mean?
Choices: [(a) highly debatable, (b) quite rare, (c) firmly established]
Answer index: 2
Subject: Language Arts
Skill: analyze-passages-from-the-outsiders-part-2
Description: Review the passage.
@@
Tim Shepard and company were already waiting when we arrived at the vacant lot, along with a gang from Brumly, one of the suburbs. Tim was a lean, catlike eighteen-year-old who looked like the model JD you see in movies and magazines. He had the right curly black hair, smoldering dark eyes, and a long scar from temple to chin where a tramp had belted him with a broken pop bottle. He had a tough, hard look to him, and his nose had been broken twice. Like Dally’s, his smile was grim and bitter. He was one of those who enjoy being a hood. The rest of his bunch were the same way. The boys from Brumly, too. Young hoods—who would grow up to be old hoods. I’d never thought about it before, but they’d just get worse as they got older, not better. I looked at Darry. He wasn’t going to be any hood when he got old. He was going to get somewhere. Living the way we do would only make him more determined to get somewhere. That’s why he’s better than the rest of us, I thought. He’s going somewhere. And I was going to be like him. I wasn’t going to live in a lousy neighborhood all my life.
From S. E. Hinton, The Outsiders. Copyright 1967 by S. E. Hinton
Look at the text in **bold**. What does this figurative language suggest?
Choices: [(a) Tim fights for survival every day., (b) Tim has always been cruel to Ponyboy., (c) Tim often does not have enough to eat., (d) Tim attempts to stay hidden at all costs.]
Answer index: 0